OpenStudy (anonymous):

How do you solve this problem 7√3- 4√6+√48- √54 ?

OpenStudy (eyust707):

There isn't anything to solve for but we can simplify it.

OpenStudy (eyust707):

There's not much we can do with \(\sqrt{3}\) or \(\sqrt{6}\) but lets take a look at \(\sqrt{48}\)

OpenStudy (eyust707):

What times what equals 48?

OpenStudy (anonymous):

6*8

OpenStudy (anonymous):

@eyust707?

OpenStudy (anonymous):

\[ 7\sqrt{3} - 4\sqrt{6} + \sqrt{48} - \sqrt{54}\] \[ 7\sqrt{3} - 4\sqrt{6} + 4\sqrt{3} - 3\sqrt{6}\] Now organize \[ 7\sqrt{3} + 4\sqrt{3} - 4\sqrt{6} - 3\sqrt{6}\] \[ 11\sqrt{3} -7\sqrt{6}\] Since we can't go any farther, below is our answer. \[ 11\sqrt{3} -7\sqrt{6}\]

OpenStudy (anonymous):

If you need more information how I got this let me know. I will explain